Please help me solve the following problems:
1. If 2-3x=11, evaluate x2(x-squared)+2x-3
2. -2[4-(3b+2)]=5-2(3b+6)
Thank you.

Answer by sdmmadam@yahoo.com(530)

1. If 2-3x=11, evaluate x2(x-squared)+2x-3
2-3x=11 ----(1)
2-11= 3x (grouping like terms,changing sign while changing side)
-9 = 3x
dividing by 3
-9/3 =x
-3 =x
That is x =-3
Putting x =-3 in x^2+2x-3
x^2+2x-3 = (-3)^2 +2X(-3)-3 = 9-6-3 = 0
Answer: Given 2-3x=11, the value of x = (-3)
And for this value of x, we get x^2+2x-3 =0
2. -2[4-(3b+2)]=5-2(3b+6)
-2X[4-3b-2]=5 -6b-12 (expanding)
-2X(2-3b)=-7-6b (grouping like terms and adding)
-4+6b = -7 -6b (expanding)
6b+6b = -7+4 (grouping like terms,changing sign while changing side)
12b = - 3
b =-3/12 = -1/4
Answer: b =(-1/4)